# CGRA4ML: A Framework to Implement Modern Neural Networks for Scientific Edge Computing

### v2

@@ II-E ML to FPGA/ASIC Frameworks
There are a few end-to-end frameworks available for hardware implementation of neural networks [12]. DNNBuilder [39] and finn [28], implement a given model as a pipeline of layers, similar to hls4ml. DNNBuilder allows the users to customize two kinds of reuse factors: channel and kernel. finn takes in an ONNX model, possibly exported from Brevitas [40], to generate Vivado HLS IP. This can be verified in simulation before implementing on AMD Xilinx FPGAs. FINN provides a PYNQ driver for prototyping.
@@
AMD Xilinx Vitis AI [29] is a closed-source library that implements a Deep Learning Processing Unit - an 8-bit micro-coded processor to process neural networks optimized through their stack on AMD Xilinx FPGAs. OpenVino is a similar stack for Intel FPGAs. Apache TVM, an open-source framework for embedded AI, implements a Versatile Tensor Accelerator (VTA) [27] as a GEMM processor using Xilinx HLS. LeFlow [41] emits HLS code, which has similar advantages and disadvantages as hls4ml.
@@
While each of these frameworks targets different groups of users, the limitations in their approaches make them incompatible with the requirements of the scientific computing community, as listed in Table I. The popularity of hls4ml and its features: arbitrary quantization, Xilinx and Intel FPGA backends, and support for limited verification, demonstrate the unique needs of the scientific edge community. The models used in scientific applications require quantization to arbitrary bit-widths, which is made possible by qkeras, and not supported by Vitis AI, OpenVino, or CGRA-ME, among others. In addition, their workflow involves implementing their models on FPGAs and later moving to ASIC designs, which is not possible with Vitis AI, Apache VTA, and OpenVino since their backends are implemented in vendor-specific HLS. CGRA4ML fills this gap by providing a high-level API where the users can first configure a CGRA and export it as SystemVerilog RTL for ASIC or FPGA implementation from any vendor, then build models quantized to arbitrary bitwidths, train them using qkeras, and export the model runtime to be executed in any host CPU.
III CGRA4ML Overview
CGRA4ML is designed to satisfy the unique requirements of the scientific edge computing community, borrowing from the distinctive strengths of hls4ml and improving upon its limitations. This section describes the CGRA4ML workflow, as shown in Fig. 1 for implementing scientific computing applications, while Sec. V explains the technical details.
A typical scientific computing workflow involves building custom models with quantization-aware training. CGRA4ML’s Python API supports this workflow in the same way as hls4ml. qkeras performs quantization-aware training. qkeras is a library that offers robust quantization options and training algorithms [14]. After training, the user exports the model, which generates intermediate inputs and outputs as integer versions of the fixed-point representation for testing.
After defining the QKeras model, it is lowered into a list of bundles. A bundle is subclass of a qkeras layer that exists in the CGRA4ML library. Its C runtime defines its functionality and runtime parameters that map the QKeras layer to the CGRA. Bundles are further described in Sec. V-A.
CGRA4ML implements large and diverse neural networks by creating an efficient yet programmable computational array that can easily be shared across multiple neural network layers. The array’s computation and data movement are design-time parameterizable (Sec. IV-A) and run-time programmable (Sec. IV-B) to accommodate a variety of DNN layer types. CGRA4ML focuses on flexibility and data movement as a key enabler for efficient scientific neural network hardware acceleration. The CGRA and the system around it is implemented as SystemVerilog RTL, to target any FPGA or ASIC flow. This fundamentally differs from hls4ml, which emits HLS code that requires a separate HLS tool to lower its output to an RTL description.
CGRA4ML provides toolflows for ASIC implementation using Cadence and Synopsys tools. Users can add paths for their technology PDKs and memory compilers to the scripts and invoke them to synthesize place and route, check DRC violations, verify layout vs. schematic (LVS), and export GDSII files. CGRA4ML allows netlist verification using gate-level simulation. Users get power, performance, and area estimates through the generated reports. ASIC designers can edit and fine-tune the base scripts generated by CGRA4ML before the final tape-out. A system on chip generated through this workflow is being taped out with the ARM-based open-source NanoSoC platform as a proof of concept, as described in Sec. V-E2.
After the CGRA4ML design is implemented, it must be integrated into an SoC. CGRA4ML generates an AXI-compliant [42] high-performance IP powered by built-in open-source, design tool-agnostic AXI DMAs that can be integrated into different SoCs. CGRA4ML also directly generates the C firmware. The firmware is designed as a simple header-only library, allowing users to get the design working within minutes and then extend it by building the system-on-chip around it, adding other hardware for co-processing or moving data through external memory, ethernet, PCIe, etc. The firmware configures the CGRA and receives data to post-process and store in buffers for processing by the next layer. The workload is shared between the CGRA and the host CPU as described in Section V-C.
Verification dominates design time [43]. Thus, ensuring that the CGRA4ML implementation matches the neural network model is critical. CGRA4ML verification infrastructure can be invoked from Python API to target Verilator, AMD Xilinx Xsim, or Synopsys VCS simulators. This verifies the user’s neural network model, the SystemVerilog RTL of the generated CGRA hardware, and the generated C runtime together through a randomized SystemVerilog testbench and reports the error between the outputs of the user-defined Python model and the generated hardware.
IV CGRA4ML Hardware Architecture
CGRA4ML is a parametrizable, efficient CGRA optimized to execute complex neural network models. CGRA4ML uses on-chip memory to cache and reuse weights, reducing data movement and energy, a pixel shifter to facilitate data reuse for intermediate results, three high-performance, open-source AXI DMAs that move data in and out of the engine, and a hardware-based DMA controller with an AXI-Lite configuration. The engine is described as a vendor-agnostic SystemVerilog design. CGRA4ML generates the C firmware to control the engine and perform data post-processing. Figure 3 gives an overview of the CGRA architecture, interfaces, and firmware.
@@ IV CGRA4ML Hardware Architecture
Figure 3: CGRA4ML architecture generated for AMD Xilinx Zynq SoC: Our CGRA-based engine (a) features one AXI-Lite port for configuration and three 128-bit AXI Manager ports for data movement. The CGRA engine (b) performs compute-heavy tasks. The engine has parameterizable bit widths, weights cache, and processing elements (PE) to fit the requirements of different models and meet hardware resource and performance constraints. CGRA4ML emits software for the ARM processor (c) to perform pixel-wise operations and handle the edge cases. .
The CGRA4ML engine features three high-bandwidth AXI-4 manager ports (up to 128-bit wide) to move inputs, weights, and outputs, and one AXI-Lite subordinate port to configure the engine with runtime parameters (see Table  II). Three AXI DMAs operate asynchronously to pull input and weight data from off-chip memory, feed them as AXI streams, receive the CGRA outputs as an AXI stream, and store them in on-chip memory. The control generator provides the configuration bits to the weights cache and pixel shifter, which process the configuration into a few bits and append them to the AXI-Stream data moving through the CGRA. CGRA4ML uses the industry-standard AMBA AXI interfaces, which enhance portability and ease the integration of the CGRA into the SoC.
IV-A Parameterized CGRA Engine Definition and Generation
A user sets the CGRA engine parameters using the Python API, which includes attributes of CGRA specification:
  • Number of Rows & Columns of PEs in the CGRA
@@ IV-A Parameterized CGRA Engine Definition and Generation
  • Bitwidth of inputs, kernel, outputs, and bias
@@
  • Frequency, I/O delays
  • Max. batch, kernel sizes, in-channels, height, and width
  • Valid and Ready probabilities for randomized verification
@@
The specification defines the CGRA architecture, which can be reprogrammed at runtime within these constraints to execute many different models. These attributes are parameterized into the SystemVerilog hardware to generate different CGRA engines quickly. The parameters are made static before synthesis based on the user-defined attributes.
The rows and columns dictate the number of PEs in the CGRA, which defines the hardware resource usage and the peak performance. The depth of the weights cache controls the weight reuse factor. These attributes provide a trade-off between performance and available hardware resources. While any model can be run on any specification, the hardware parameters can be optimized for specific models as described in Subsection IV-C. The bitwidths are customizable across the inputs, outputs, kernel, and bias. The PE datapaths are automatically adapted to match the requirements of these data types. Quantization is critical to meet high throughput and low latency requirements in scientific applications; thus, CGRA4ML supports fixed-point datatypes. The AXI interfaces are parameterized to easily match the SoC interfaces. The maximum limits of layer dimensions determine the register sizes within the hardware (see Table II).
CGRA4ML provides a robust verification environment using the randomized transactional testbench suite. The valid and ready probabilities define the randomness in toggling the s_valid and m_ready signals of the AXI interfaces. Setting the probabilities to 1 gives the fastest simulation, where data is transferred every cycle. This is helpful as a smoke test when a user intends to modify the hardware and debug their modification. Setting it to lower values, e.g., 0.1 and 0.01, stress tests the hardware using randomized transactions, exposing hardware bugs that smoke tests cannot reveal.
@@
Fig. 3 (d) shows the CGRA4ML PEs, which include a multiplier, accumulator, two registers, and two multiplexers. Simplicity is favored over complexity to increase computational density and efficiency. CGRAs designed with a higher degree of flexibility in data movement and operation types suffer from the complexity required to program them and often utilize their resources less effectively. Therefore, our lightweight CGRA design aims to offer minimal programming complexity while supporting high utilization required for scientific edge computing. The PE columns can group themselves and process convolution kernels, sharing data to maximize data reuse. The pixel shifter maximizes data reuse across vertical convolution while maintaining modularity. The weights cache consists of two on-chip RAMs with parameterized depths in a ping-pong configuration to allow full throughput.
TABLE II: Runtime parameters of the unified dataflow

### v3

@@ 2.5. ML to FPGA/ASIC Frontend Frameworks
There are many end-to-end frameworks available for hardware implementation of neural networks (Rahimifar et al., 2022). DNNBuilder (Zhang et al., 2018) and finn (Umuroglu et al., 2017), implement a given model as a pipeline of layers, similar to hls4ml. DNNBuilder allows the users to customize two kinds of reuse factors: channel and kernel. finn takes in an ONNX model, possibly exported from Brevitas (Pappalardo, 2024), to generate Vivado HLS IP. This can be verified in simulation before implementing on AMD Xilinx FPGAs. FINN provides a PYNQ driver for prototyping.
@@
AMD Xilinx Vitis AI (AMD-Xilinx, 2024) is a closed-source library that implements a Deep Learning Processing Unit, an 8-bit micro-coded processor to process neural networks optimized through their stack on AMD Xilinx FPGAs. OpenVino is a similar stack for Intel FPGAs. Apache TVM, an open-source framework for embedded AI, implements a Versatile Tensor Accelerator (VTA) (Moreau et al., 2018) as a GEMM processor using Xilinx HLS. LeFlow (Noronha et al., 2018) emits HLS code, which has similar advantages and disadvantages as hls4ml.
Recent compiler stacks such as PyTorch Dynamo (Team, 2025b) and OpenAI Triton (Team, 2025a) focus on automatic kernel fusion and code generation for GPUs/CPUs (with early explorations for other backends). Our contribution is orthogonal: cgra4ml targets a CGRA-style spatial accelerator with unified dataflow and a firmware/runtime that binds those schedules to a memory-mapped, DMA-driven engine on FPGA/ASIC. In principle, TorchInductor/Triton-style schedules could be lowered to our ISA/dataflow with a translation layer; we leave such integration to future work.
@@
While each of these frameworks targets different kinds of users, the limitations in their approaches make them incompatible with the requirements of the scientific computing community, as listed in Table 1. The popularity of hls4ml and its features: arbitrary quantization, Xilinx and Intel FPGA backends, and support for limited verification, demonstrate the unique needs of the scientific edge community. The models used in scientific applications require quantization to arbitrary bit-widths, which is made possible by qkeras, and not supported by Vitis AI, OpenVino, or CGRA-ME, among others. In addition, their workflow involves implementing their models on FPGAs and later moving to ASIC designs, which is not possible with Vitis AI, Apache VTA, and OpenVino since their backends are implemented in vendor-specific HLS.
2.6. Motivation for cgra4ml
Modern scientific applications increasingly rely on neural networks to process data rapidly and efficiently at the edge, imposing stringent requirements on throughput, latency, and power efficiency. hls4ml is a popular framework among scientists—deployed in custom ASICs for L1 triggers and on FPGAs in L2 triggers at the LHC, among other uses—because it offers an accessible, dataflow-style, layer-by-layer implementation for small networks and a range of precisions. However, this approach scales poorly beyond a few layers: on-chip buffering and per-layer specialization grow quickly, making larger models difficult to fit on FPGAs and small ASICs. Reusable accelerators amortize resources across layers and are therefore more suitable for larger networks, yet most available designs provide limited, non–user-friendly front ends and often lack features needed by the scientific computing community, including sub–8-bit quantization, multiple backends (FPGA and ASIC), end-to-end verification, and runtime generation. Moreover, hls4ml and other HLS-based DNN frameworks are tied to vendor-specific toolchains, which constrains backend portability and demands substantial engineering to build complete systems with a host CPU, DMA subsystems, drivers, and production firmware.
As summarized in Fig. 3, hls4ml remains the community’s go-to for dataflow implementations over varying bitwidths and shallow-to-moderate depths, while finn likewise follows a dataflow paradigm and excels at ultra–low-precision designs. Traditional AI accelerators, in contrast, target large models with $\geq$8-bit quantization. cgra4ml is designed to fill the gap between these regimes by enabling models with sub–8-bit quantization that are too large to realize with hls4ml or finn, while providing multiple backends, end-to-end verification, and automatic runtime/driver generation to streamline full-system integration.
3. cgra4ml Overview
The needs of the scientific computing community for a framework that can implement high-performance and programmable accelerators using the familiar hls4ml-style interface, with the generation of cross-platform hardware and production-ready firmware, were identified in Section 2. Often models exceed on-chip memory capacities, demanding efficient off-chip data handling and more flexible hardware configurations. And typically users would default to not using larger models due to the inability for hls4ml to implement them. To address these limitations, we introduce cgra4ml, an open-source, modular framework explicitly designed for the scientific edge computing community that can better facilitate reuse across layers, implement larger models, and ease integration into larger systems. cgra4ml leverages the strengths of hls4ml: ease of use, quantization-aware training, and rapid hardware accelerated deployment, while substantially extending capabilities in several key areas.
Firstly, cgra4ml supports large-scale neural network implementations by employing a parametrizable CGRA, enabling efficient spatial reuse of processing elements and optimized off-chip data movements. cgra4ml outputs SystemVerilog RTL eliminating the need for HLS tools required by other frameworks. This RTL-centric approach streamlines FPGA implementation and facilitates seamless ASIC integration, thereby providing flexibility across different hardware targets without extensive code modifications. Furthermore, cgra4ml includes an extensive verification framework integrated within its workflow, significantly reducing the time spent in debugging and validation processes, which are notoriously time-consuming in hardware design. This section describes the user workflow of cgra4ml and the Python frontend that makes our infrastructure possible.
Refer to caption cgra4ml workflow
Figure 4. cgra4ml workflow as outlined in Sec. 3. Users first build quantized neural networks and train them in a quantization-aware manner using qkeras (Sec. 4.1). Users then define a CGRA definition, which generates vendor-agnostic SystemVerilog RTL hardware specifications and TCL tool flows (Sec. 4.1). The model can then be exported to generate weights and a C runtime firmware (Sec. 4.2.2). The generated hardware IP is then verified comprehensively with the model and firmware (Sec. 10) using our randomized, transactional SystemVerilog testbench suite with DPI-C extensions. Finally, the bitstream generated from the FPGA toolflow and the C firmware can be loaded into an FPGA to be tested in seconds (Sec. 5.3). After such rapid prototyping, the same hardware design can be moved to ASIC (Sec. 5.4).
The overarching goal of cgra4ml is to empower scientists and hardware designers by abstracting away intricate hardware optimization details that slow down development for non-expert users. cgra4ml’s user workflow is intuitive, streamlined, and flexible, supporting rapid development cycles and iterative experimentation. Users define neural network architectures via a Python API leveraging qkeras for quantization-aware training, after which cgra4ml automatically generates hardware specifications, runtime firmware, and robust verification suites. This end-to-end approach enables rapid prototyping, flexible experimentation, and efficient deployment, ultimately allowing users to prioritize innovative applications over hardware intricacies. It comprises the following steps:
3.0.1. Model Definition and Training
Users begin by defining neural network models using the Python-based qkeras library, which supports quantization-aware training. This step ensures models are optimized for low-precision hardware implementations while maintaining accuracy.
3.0.2. Accelerator Configuration
After training, users specify static parameters for the CGRA architecture, including the number of processing elements (PEs), bitwidths, and memory hierarchies. Users can adjust these parameters to tune the design to within the resource constraints of the target hardware.
3.0.3. RTL Generation
Using the specified parameters, cgra4ml automatically generates vendor-agnostic SystemVerilog RTL code. This eliminates dependency on vendor-specific tools and simplifies porting between FPGA and ASIC targets.
3.0.4. Firmware Generation
cgra4ml produces optimized runtime firmware, facilitating seamless integration and operation of the generated hardware within an SoC environment. Users can quickly deploy and test their designs in real-world applications.
3.0.5. Verification
The generated hardware and firmware undergo comprehensive verification using cgra4ml ’s built-in randomized transactional testbench suite. This step ensures functional correctness and reliability prior to physical deployment.
3.0.6. FPGA/ASIC Deployment
Finally, users deploy their verified designs onto FPGA platforms for rapid prototyping and experimentation. Once validated, the same RTL design seamlessly transitions to ASIC flows.
Through this structured, hls4ml-inspired workflow, cgra4ml enables users to rapidly move from model concept to physical deployment, streamlining innovation, and accelerating scientific discovery.
4. cgra4ml: Infrastructure
To bridge the gap between high-level neural network development and hardware-accelerated edge deployment, cgra4ml provides a modular, holistic infrastructure that spans model quantization, hardware generation, firmware synthesis, and system-level verification. Central to this flow is the concept of bundles, which are modular, deterministic units of execution that encapsulate groups of DNN layers, making them suitable for hardware acceleration. These bundles allow for clean partitioning of compute workloads between a coarse-grained reconfigurable array (CGRA) and a general-purpose CPU. This design philosophy supports scalability, reuse, and extensibility, enabling developers to experiment with diverse model structures while maintaining control over system-level performance and resource utilization. The backend pipeline, which comprises SystemVerilog CGRA RTL, portable runtime firmware, and automated toolchains for FPGA and ASIC flows, ensures that scientific edge applications can transition from simulation to physical deployment with minimal effort.
4.1. Frontend Infrastructure: Model to Bundles
Figure 5. Bundle: A given DNN is decomposed into a list of bundles, where a bundle is a group of layers that can be deterministically executed by the system generated by cgra4ml. The CGRA accelerates the simple but compute-heavy operations, while the CPU executes the complex but lightweight pixel-wise operations. This flexibility allows us to add new, complex operations easily.
A given DNN is decomposed into a list of bundles
Modern neural networks have complex, seemingly arbitrarily structured graphs of computational layers. Therefore, a framework should balance the flexibility it offers to the users to the optimized backend implementation of these networks in hardware. This requires a modular and extensible intermediate representation (IR). The frontend of cgra4ml is responsible for translating a high-level neural network model into a hardware-ready specification. This process is centered around the concept of bundles, which are modular, hardware-executable units that serve as the bridge between qkeras models and cgra4ml accelerator.
Users construct and train quantized models using qkeras, a widely adopted library from Google that enables fine-grained control over quantization levels and supports straight-through estimators for backpropagation. Once trained, cgra4ml parses the model’s computational graph and decomposes it into a list of bundles. Each bundle represents a deterministic sequence of layers that can be directly mapped onto the CGRA. In the Python frontend, a bundle is implemented as a subclass of a qkeras layer, enriched with metadata and configuration parameters required for hardware mapping. In the backend firmware, each bundle corresponds to a structured set of runtime instructions, including memory access patterns, kernel parameters, and execution schedules.
A bundle acts as the core building block of the cgra4ml infrastructure and enables hardware/software partitioning. Compute-intensive operations such as convolutions and matrix multiplications are offloaded to the CGRA, while lightweight and control-heavy operations such as pooling or normalization are executed by the CPU. Convolutional and dense layers involve large numbers of multiply-accumulate (MAC) operations and benefit most from parallelization on the CGRA. The CGRA features a parameterized 2D array of processing elements ($R{\times}C$) that execute MAC operations in parallel and store results in on-chip memory. Meanwhile, CPU-executed tasks handle edge conditions and pixel-wise computations.
Figure 5 illustrates how a ResNet-50 model is decomposed into bundles, each encompassing layers such as Conv2D, activation, pooling, flattening, and skip connections. The Python frontend transforms the model, compares intermediate results against qkeras outputs, and partitions execution between the CGRA and CPU. It selectively enables layers, and passes their configuration as runtime parameters inferred from the user defined model. The generated C firmware then orchestrates bundle execution at runtime in a way that ensures functional equivalence to the Python model.
cgra4ml also empowers users to define new bundles by extending the Python frontend and adding corresponding runtime definitions in C. These composite operations are decomposed during compilation into a static runtime, allowing users to define and deploy more complex architectures such as residual networks without modifying the core hardware design. This architectural decision is made to keep the hardware complexity minimal, and to enable new operations such as activation functions when necessary. We characterize the CPU execution latency cost in Table 4 and outline our future roadmap to move the writeback stage into hardware in Sec. 6.2.
4.2. Backend Infrastructure: CGRA RTL and Firmware Generation
Figure 6. Modular and extensible infrastructure of the cgra4ml framework as described in Section 4. The Python API built around Google’s qkeras, extracts the fixed-point representations of intermediate tensors, generating the binary weights and model specification. The Python frontend also generates vendor-agnostic, synthesizable SystemVerilog RTL design of the CGRA per user specifications, optimized C runtime firmware and TCL toolflows for ASIC and FPGA implementations. Our comprehensive verification suite powered by randomized, transactional, SystemVerilog testbenches interfacing with the C firmware via DPI-C, then verifies the user’s model, the generated CGRA, and runtime together as it would run on an embedded system, emulating cases like memory congestion. The CI/CD pipeline tests the entire framework on a set of models over different CGRA configurations
Modular and extensible infrastructure of the cgra4ml framework as described
The backend of cgra4ml translates the high-level bundle specifications into hardware and runtime components, enabling deployment across FPGA and ASIC platforms. It encompasses three primary components: RTL generation, firmware synthesis, and toolchain integration for verification and implementation. This subsection outlines these steps in the backend process, while the following subsections discuss each in detail.
4.2.1. RTL Generation
cgra4ml produces synthesizable, vendor-agnostic SystemVerilog RTL for a parametrized CGRA engine. Users configure key static parameters, such as the number of processing elements, memory hierarchy depth, and data bitwidths, via the Python frontend. These parameters are embedded into the hardware description to generate an optimized datapath and control logic. The default datatype is fixed-point. However, the users may integrate a floating-point MAC by specifying its latency of the multiplier and accumulator in our parameter list. The resulting RTL supports standard AXI interfaces for seamless SoC integration. After implementation, the design must be integrated into a broader SoC. cgra4ml generates a high-performance, AXI-compliant IP powered by open-source, design tool-agnostic AXI DMAs. These modules simplify integration into custom SoCs or evaluation platforms. The firmware and hardware together ensure efficient data transfer and workload sharing between the CGRA and the host processor.
4.2.2. Firmware Generation
Alongside RTL, cgra4ml generates modular C firmware that manages CGRA execution and coordinates software-side tensor operations. This firmware is structured in three layers: (1) a model-specific configuration header exported from Python, (2) a portable runtime layer for buffer management and synchronization, and (3) architecture-specific routines for platform integration. The firmware is designed as a simple header-only library with a thin Hardware Abstraction Layer (HAL), allowing users to quickly get the design working and extend it to include co-processors and memory interfaces like Ethernet and PCIe. This modularity allows the firmware to target diverse embedded processors including ARM, RISC-V, and x86, and to be reused across different CGRA configurations. The runtime targets a stable hardware contract via a custom bank of 32-bit registers for control and status, with DMA descriptors. No changes to the unified dataflow or runtime firmware are required when targeting different architectures and FPGA/ASIC backends. Only memory macro selection and interconnect adapters vary by process and SoC fabric. We have validated the portability of our firmware on ZYNQ Programmable SoCs with FPGAs and ARM Cortex processors, on Ibex SoC with a RISC-V processor, and on x86 CPUs for verification.
4.2.3. Toolchain Integration
cgra4ml includes fully automated toolchains for FPGA and ASIC flows. For FPGAs, it generates TCL scripts that configure IPs, integrate the CGRA into SoC shells (e.g., Zynq), and synthesize/place/route the design. For ASICs, it produces scripts for Cadence Genus and Innovus, or Synopsys DC and ICC, enabling synthesis, physical design, and GDSII export. Users can customize these flows by linking to their PDKs, memory compilers, and setting specific constraints. Reports generated through these tools provide power, performance, and area (PPA) estimates. A proof-of-concept SoC using cgra4ml IP integrated into the ARM-based NanoSoC platform demonstrates the viability of this flow.
4.2.4. System Verification
A central component of the backend is the DPI-C-enabled verification suite. Testbenches written in SystemVerilog interface with the C firmware to simulate the complete stack: from quantized model outputs to hardware response. Randomized AXI stimuli stress-test corner cases such as memory congestion or timing edge conditions. This approach reduces the verification burden and ensures fidelity between the model and its physical implementation.
Together, these backend systems allow cgra4ml to bridge the gap between machine learning model development and production-grade hardware deployment, offering flexibility, correctness, and performance portability across platforms.
4.3. Parameterized CGRA Engine Architecture
@@
Figure 7. cgra4ml architecture generated for AMD Xilinx Zynq SoC: Our CGRA-based engine (a) features one AXI-Lite port for configuration and three 128-bit AXI Manager ports for data movement. The CGRA engine (b) performs compute-heavy tasks. The engine has parameterizable bit widths, weights cache, and processing elements (PE) to fit the requirements of different models and meet hardware resource and performance constraints. cgra4ml emits software for the ARM processor (c) to perform pixel-wise operations and handle the edge cases.
cgra4ml architecture generated for AMD Xilinx Zynq SoC
cgra4ml generates a system of hardware components, as described in Fig. 7, that work together with a host processor to process the user defined neural network model. The components are designed to have AXI interfaces for modular hardware design. The system consists of the following subsystems, each with specified parameters that the user sets using the Python API:
  • CGRA - A Coarse Grained Reconfigurable Array (CGRA) with AXI-stream interfaces.
    • Number of Rows & Columns of processing elements (PEs)
    • Bitwidth of inputs, kernel, outputs, and bias
  • Weights cache - an AXI stream module with ping-pong buffers of on-chip SRAM for maximizing data reuse, counters and controllers.
@@
    • Maximum batch, kernel sizes, in-channels, height, and width
  • Three Direct Memory Access (DMA) modules that convert between AXI stream and full AXI4.
@@
  • Verification IPs and wrappers for full system Verification
    • Valid and Ready probabilities for randomization
In addition to these tunable parameters, clock frequency and I/O delays of the full system are also parametrizable.
cgra4ml implements large and diverse neural networks by creating an efficient yet programmable computational array that can easily be shared across multiple neural network layers. The array’s computation and data movement are design-time parameterizable and run-time programmable to accommodate a variety of DNN layer types. The CGRA is designed to perform various types of neural network layers with optimal performance. This subsection discusses the static, design-time parametrization of the CGRA, while the next subsection discusses runtime configuration.
@@
The static hardware specification defines the CGRA architecture, which can be reprogrammed at runtime within these constraints to execute many different models. These attributes are parameterized into the SystemVerilog hardware to generate different CGRA engines quickly. The parameters are made static before synthesis based on the user-defined attributes. cgra4ml supports fixed-point datatypes, since quantization is critical to meet high throughput and low latency requirements in scientific applications. Their bitwidths are customizable across the inputs, outputs, kernel, and bias. The PE datapaths are automatically adapted to match the requirements of these data types.
@@
The rows and columns dictate the number of PEs in the CGRA, which defines the hardware resource usage and the peak performance. Fig. 7 (d) shows the internal architecture of PEs, which include a multiplier, accumulator, two registers, and two multiplexers. Simplicity is favored over complexity to increase computational density and efficiency. CGRAs designed with a higher degree of flexibility in data movement and operation types suffer from the complexity required to program them and often utilize their resources less effectively. Therefore, our lightweight CGRA design aims to offer minimal programming complexity while supporting high utilization required for scientific edge computing. The PE columns can group themselves and process convolution kernels, sharing data to maximize data reuse.
The depth of the weights cache controls the weight reuse factor. The weights cache consists of two on-chip RAMs with parameterized depths in a ping-pong configuration to allow full throughput. The pixel shifter maximizes data reuse across vertical convolution while maintaining modularity. Their dimensions provide a trade-off between performance and available hardware resources. While any model can be run on any specification, the hardware parameters can be optimized for specific models as described in Subsection 6.1. The maximum limits of layer dimensions determine the register sizes within the hardware (see Table 2).
As shown in Fig. 7, our hardware modules feature open AXI-Full, AXI-Stream, and AXI-Lite interfaces to ensure compatibility with a wide variety of SoCs and other hardware IPs. We utilize open-source AXI and AXI-Stream cores (Alex Forencich, 2024; Alex Forencich., 2024) in our SoC generation, modified to support ASIC flow in addition to FPGA. The AXI interfaces are parameterized to easily match the SoC interfaces. All these modules are wrapped into a single IP that can be easily instantiated and connected to any system with AXI ports.
4.4. Runtime Dynamic Reconfiguration and Dataflow
Table 2. Runtime parameters of the unified dataflow